Part A: Suppose n1=19, n2=21 and n3=23. For a one-factor ANOVA, the dferror would be __________....

Part A:

Suppose n1=19, n2=21 and n3=23. For a one-factor ANOVA, the dferror would be __________.

Part B:

Suppose n1=19, n2=21 and n3=23. For a one-factor ANOVA, the dftreatwould be ______.
Answer #1

We have here,

k=3,

and

n=19+21+23=63

Part A

the dferror would be = n-k=63-3=60

Part B

the dftreatwould be = k - 1 = 3 - 1 = 2